1. The following three forces act on an object, determine the magnitude and direction of the resultant force:     (4 mks)
  •  8N due West
  • 4N due South
  • 5N due East


  1. A box resting on the table has a downward force of 500 N and an upward force of 500 N. 

What is the resultant force on the box? (1 mk)

What conclusion can we make about the box? (1 mk)


The following four forces are acting on an object, find the magnitude and resultant force acting on the object: (4 mks)

  • 150N due East
  • 950N due South
  • 500N due West
  • 750N due North
1
Expert's answer
2021-03-14T19:14:17-0400

1)


"F=\sqrt{(8-5)^2+4^2}=5\ N\\
\theta=\arctan{\frac{3}{4}}=37\degree \text{ W of S}"

2)


"F=500-500=0\ N"

No motion .

3)


"F=\sqrt{(500-150)^2+(850-750)^2}=364\ N\\
\theta=\arctan{\frac{100}{350}}=16\degree \text{ S of W}"